OpenStudy (anonymous):

In studying the gene pool of a population, you find that 21% of the population exhibits the recessive trait for eye color and the other 79% of the population exhibits the dominant trait. The dominant allele has complete dominance over the recessive allele. How would you use this information to solve for the frequency of the recessive allele?

OpenStudy (anonymous):

I think 12

OpenStudy (anonymous):

does that help

OpenStudy (aaronq):

The only equations you need for these types of problems: p + q =1 p = frequency of the dominant allele in the population q = frequency of the recessive allele in the population and p^2 + 2pq + q^2 =1 p^2 = percentage of homozygous dominant individuals q^2 = percentage of homozygous recessive individuals 2pq = percentage of heterozygous individuals